A certain radioactive substance has a half-life of 400 years. If 60 kg is accidently released when will 0.25 kg remain? (assumming it decays continuously)
Please help me with this one, I'm not even sure where to start! Don't you just love the assignments we get before we actually learn the material!! I hate to ask for help, but I have no idea how to set this up and calculate it!
Thank you,
Tracy

Answer by scott8148(6628)

the general formula for half-life decay is ___ A = Ao [.5^(t / h)]

A = amount after decay
Ao = starting amount
t = time
h = half-life

ln(A / Ao) = (t / h)[ln(.5)] ___ h[ln(A / Ao)] / [ln(.5)] = t